# Work-From-Home and Privacy: What Do Workers Face and What are They Doing About it?

### v1

@@ Abstract.
The COVID-19 pandemic has reshaped the way people work, normalizing the practice of working from home (WFH). However, WFH can cause a blurring of personal and professional boundaries, surfacing new privacy issues, especially when workers take work meetings from their homes. As WFH arrangements are now standard practice in many organizations, addressing the associated privacy concerns should be a key part of creating healthy work environments for workers. To this end, we conducted a scenario-based survey with 214 US-based workers who currently work from home regularly. Our results suggest that privacy invasions are commonly experienced while working from home and cause discomfort to many workers. However, only a minority said that the discomfort escalated to cause harm to them or others, and the harm was almost always psychological. While scenarios that restrict worker autonomy (prohibit turning off camera or microphone) are the least experienced scenarios, they are associated with the highest reported discomfort. In addition, participants reported measures that violated or would violate their employer’s autonomy-restricting rules to protect their privacy. We also find that conference tool settings that can prevent privacy invasions are not widely used compared to manual privacy-protective measures. Our findings provide better understanding of the privacy challenges landscape that WFH workers face and how they address them. Furthermore, our discussion raised open questions that can inspire future work.
Work-from-home (WFH), is a flexible work setting that allows workers to perform their work duties, including collaborative work with others such as meeting and teaching, remotely from their homes instead of a dedicated office space provided by employers. WFH often requires the use of information and communication technologies, such as voice and video conferencing tools (e.g. Zoom and Teams). Before 2020, WFH was often provided as a benefit that gives workers more flexibility (Chiru, 2017; Hayes et al., 2021; Ford et al., 2021). However, in early 2020, with the advent of the global Coronavirus (COVID-19) pandemic (World Health Organization, 2024), most workers who used to work from traditional offices found themselves required (by their employers or jurisdiction laws) to adhere to lockdown rules and work from home to prevent the virus spread. WFH was no longer a benefit, but a mandate, despite the challenges it poses to some workers (Ralph et al., 2020; Ford et al., 2021). WFH can be especially challenging for workers who live with children, lack appropriate space at home, or lack the skills needed to deal with new technologies. Some WFH challenges were exemplified in viral videos. For example, in 2017 Professor Robert Kelly’s two young children invaded a live-streaming interview about South Korea’s politics on the British Broadcasting Channel (BBC) and were subsequently dragged out of the room by his wife (Chappell, 2020), as illustrated in Figure 1.
@@
Post-COVID-19, many individuals and organizations find that the benefits of WFH outweigh the drawbacks. Many organizations have changed their policies to allow WFH on an ongoing basis (Smite et al., 2023; Christian, 2023; Lufkin, 2022). According to a 2024 report by WFHResearch on US workers, the growth in WFH post-COVID-19 is nearly equivalent to 40 years of pre-COVID-19 growth (Barrero et al., 2024). The same report shows that as of early 2024, 13% of full-time workers were working fully remotely, and 29% were in a hybrid mode (i.e., some days from home, while other days from office space provided by employers). Data by the US Bureau of Labor Statistics shows that as of February 2024, 22.7% of workers work from home at least some of the time (US Bureau of Labor Statistics, 2023).
With a significant fraction of the American workforce working from home all or part of the time, addressing the privacy concerns and behaviors that first arose out of emergency WFH arrangements should now be considered a key part of creating healthy work environments for workers. To do that, we must first understand the privacy concerns and behaviors of those whose “new normal” includes WFH on a regular basis.
To this end, we surveyed 214 US-based workers who currently work from home regularly about their experiences with privacy-invasive scenarios pertaining to: audio, video, data [^1] For brevity, throughout the paper we use the word “data” to denote “computer/browsing data.”, and autonomy, that can occur in a WFH setting. We asked them whether they experienced the scenarios in the past three years, and if so, how comfortable or uncomfortable they felt. We also asked them whether they had taken any measures to prevent the scenarios, what those measures were, and if not, why they had not taken any preventative measures. To gain more insights into the characteristics of privacy-invasive scenarios that cause discomfort, our survey asked the participants follow-up questions about the most memorable scenario that made them feel uncomfortable.
@@
    Privacy-invasive scenarios are commonly experienced and caused discomfort to most of our participants. Nearly all participants (93.9%) experienced at least one of the privacy-invasive scenarios, and over half of participants (65.4%) felt uncomfortable during at least one of the scenarios they experienced.
@@
    While scenarios that restrict worker autonomy such as prohibiting them from turning off camera or microphone are not among the most commonly experienced scenarios, they are associated with the highest reported discomfort. Participants often violated or would violate employer’s autonomy-restricting rules to protect their privacy.
@@
    Participants reported experiencing more discomfort with audio privacy invasions than with video privacy invasions (46.2% of audio privacy-invasive scenarios caused discomfort vs. 28.8% of video privacy-invasive scenarios).
@@
    Smart measures (software/hardware settings or features to prevent privacy-invasive scenarios, e.g. virtual background and noise cancellation features) are not widely adopted compared to manual measures (e.g. manually covering the camera and turning on/off the microphone). When used, more participants reported using settings or features to prevent video scenarios (24.0%) than audio (13.11%), data (8.4%), and autonomy (4.2%) scenarios.
@@
    Most participants who experienced a privacy-invasive scenario that made them feel uncomfortable said that the incident did not escalate to cause harm to them or others (74.3%).
Our findings contribute to our understanding of the privacy challenges that WFH workers face and how they address them, which can help in identifying ways to improve WFH experiences. Furthermore, we provide several open questions that inspire future work.
1. Related Work
As Cho et al. noted, for many workers, the shift to WFH during the pandemic was not optional and many workers were “forced” to work from home, even though their homes might be inconvenient for non-domestic purposes such as work or study (Cho et al., 2022).
Emami-Naeini et al. conducted a survey on privacy concerns and attitudes toward remote communications (Emami-Naeini et al., 2021). They found that privacy and security issues in remote communications are among the issues that people care about, and affect their comfort in using conferencing tools (Emami-Naeini et al., 2021). They also found that most participants in all remote communication contexts (work, socializing, and learning) felt that they lacked autonomy in choosing the conferencing tool and/or using the microphone and camera in their remote communications (Emami-Naeini et al., 2021). Prange et al. investigated privacy breaches that have occurred as a result of video-based meetings (Prange et al., 2022). This research was not focused solely on WFH and participants shared experiences relating to both personal and business-related purposes (Prange et al., 2022). Of the numerous examples of audio and video-based privacy breaches reported by their participants, over half of them occurred in business-related scenarios, and many related to video conferencing tools broadcasting accidental audio or video sharing of private situations (Prange et al., 2022). Cohney et al. surveyed university instructors and administrators, and analyzed a set of most popular virtual learning platforms (Cohney et al., 2021). Their instructor survey showed that nearly half of their participants have security and privacy concerns and care about recording controls, and some were dissatisfied with the choice of the platform and its configuration by their university (Cohney et al., 2021).
Bispham et al. interviewed six cybersecurity experts on the cybersecurity issues organizations face in the WFH setting during COVID-19 (Bispham et al., 2022). They suggested that while many cybersecurity problems such as phishing attacks have increased, many of the interviewed experts indicated the successful management of cybersecurity and that cybersecurity issues did not undermine effective WFH (Bispham et al., 2022). They also suggested the need for surveys to gain empirical perspective on the cybersecurity issues in the WFH setting, which is the contribution of our work (Bispham et al., 2022). Kotak et al. evaluated the various technologies that allow for secure remote work, and introduced a taxonomy of the security threats that these technologies and WFH introduce to organizations (Kotak et al., 2023). They identified several mitigation techniques (Kotak et al., 2023). However, while these mitigations can help improve security, they are susceptible to numerous existing and emerging threats (Kotak et al., 2023). Moreover, this research focused on the protection of organizational data and assets, as opposed to the protection of the worker’s privacy and security (Kotak et al., 2023).
@@ 1. Related Work
Hilgefort et al. revealed that software options that can be used to enhance privacy while using video conferencing tools, in particular, virtual backgrounds (which can be used to obfuscate surroundings) are also vulnerable to attack (Hilgefort et al., 2021). They found that artificial intelligence can be used to effectively remove the blurred virtual backgrounds native to many video conferencing tools, revealing the participants’ true physical setting (Hilgefort et al., 2021).
While the aforementioned studies have looked at privacy and security issues in remote communication, none of them deeply explored privacy issues while working from home, which is our focus.
@@ 2. Method
We conducted a 214-participant online survey. We obtained ethical approval for the study from the Anonymous Institution Institutional Review Board (IRB). All participants were presented with an online consent form at the beginning of the survey.
@@ 2.1. Design
Our survey consisted of two parts: a screening survey and a main survey. We used the screening survey to select participants who were currently employed (at the time of the study) and working from home regularly (one or more days per week, either in full or hybrid mode), in addition to other screening criteria. Only those who met our screening criteria were invited to take the main survey and were immediately redirected to the main survey if they agreed. The screening survey is provided in Appendix B.2.
To design the survey questions, three members of the research team developed a set of 14 privacy-invasive scenarios that can occur in a WFH setting, listed in Table 1. Some of these scenarios were inspired by the results of an IRB-approved unpublished 50-participant exploratory survey on WFH privacy behaviors and concerns conducted by some of this paper’s authors and others in 2023. The 14 scenarios were divided into four categories: audio, video, data, and autonomy. The audio and video categories (five scenarios for each category, totaling 10 scenarios) presented leaked audio/video scenarios through the subject’s microphone/camera from five different actors: the subject, another adult, child, pet, and object (e.g. appliances, books, furniture, artwork). The data category (two scenarios) presented leaked data scenarios through the subject’s shared screen. Finally, the autonomy category (two scenarios) presented restricted subject’s autonomy to turn off the microphone/camera. All scenarios were presented in the context of a work-related remote call or meeting from home.
Table 1. The 14 scenarios presented to participants were divided into categories. Scenarios that share a large amount of text are listed once in the shaded rows for each scenario category, starting with three dots (…) that represent the variable part of the scenario that changes according to the individual scenario listed below. The * denotes “Computer/Browsing.”
@@
To further characterize WFH scenarios that cause discomfort, after all scenario questions were answered, we presented participants with a list of all the scenarios that they indicated that they made them feel uncomfortable. Then, we asked participants to select the most memorable scenario, and we then asked them follow-up questions about the one selected scenario. The follow-up questions included an open-ended question to describe the incident in more detail, and several multiple-choice questions including whether the discomfort escalated to cause harm to themselves or others. If the participant did not report any scenario that made them feel uncomfortable, we asked them to describe a scenario that would make them feel uncomfortable if they were to experience it while working from home, and why. The survey concluded with demographic questions.
To avoid security and privacy priming, we did not mention the word privacy or security to the participants at any point prior to or during the survey. The survey was conducted using Qualtrics (Qualtrics, 2024), an online survey platform. The full screening and main survey questions are included in Appendix B

### v3

# Work From Home and Privacy Challenges: What Do Workers Face and What are They Doing About it?

@@ Abstract.
The COVID-19 pandemic has reshaped the way people work, normalizing the practice of working from home. However, work from home (WFH) can cause a blurring of personal and professional boundaries, surfacing new privacy issues, especially when workers take work meetings from their homes. As WFH arrangements are now standard practice in many organizations, addressing the associated privacy concerns should be a key part of creating healthy work environments for workers. To this end, we conducted a scenario-based survey with 214 US-based workers who currently work from home regularly. Our results suggest that privacy invasions are commonly experienced while working from home and cause discomfort to many workers. However, only a minority said that the discomfort escalated to cause harm to them or others and that the harm was almost always minor and psychological. While scenarios that restrict worker autonomy (prohibit turning off camera or microphone) are the least experienced scenarios, they are associated with the highest reported discomfort. In addition, participants reported measures that violated or would violate their employer’s autonomy-restricting rules to protect their privacy. We also find that conference tool settings that can prevent privacy invasions are not widely used compared to manual privacy-protective measures. Our findings provide a better understanding of the privacy challenges landscape that WFH workers face and how they address them, providing useful insights to organizations’ policymakers and technology designers for areas of improvements, to provide healthier work environments to workers.
1. Introduction
Work from home (WFH), is a flexible setting that allows workers to perform their work duties including collaboration with others such as meeting and teaching, remotely from their homes instead of a dedicated office space provided by employers. WFH often requires the use of information and communication technologies such as voice and video conferencing tools (e.g. Zoom (Inc., 2025) and Microsoft Teams (Microsoft, 2025)). Before 2020, WFH was often provided as a benefit; however, in early 2020 with the advent of the global Coronavirus pandemic (COVID-19) (World Health Organization, 2024), WFH was no longer a benefit, but a mandate, despite the challenges it poses to some workers (Ford et al., 2021; Fukumura et al., 2021; Varner and Schmidt, 2022). WFH can be especially challenging for workers who live with children, lack appropriate space at home, or lack the skills needed to deal with new technologies. Some WFH challenges were exemplified in viral videos. For example, in 2017, Professor Robert Kelly’s two young children invaded a live stream interview about South Korea’s politics on the British Broadcasting Channel (BBC) and were subsequently dragged out of the room by his wife (Chappell, 2020), as illustrated in Figure 1.
@@ 1. Introduction
After COVID-19, many organizations have changed their policies to allow WFH on an ongoing basis (Smite et al., 2023; Christian, 2023; Lufkin, 2022). According to a 2024 report by WFHResearch on US workers, the growth in WFH after COVID-19 is nearly equivalent to 40 years of growth before COVID-19 (Barrero et al., 2024). The same report shows that by early 2024, 13% of full-time workers were working fully remotely, and 29% were in a hybrid mode (Barrero et al., 2024). Data from the US Bureau of Labor Statistics show that in December 2024, 23.1% of workers worked from home at least some of the time (US Bureau of Labor Statistics, 2023).
With a significant fraction of the American workforce working from home all or part of the time, understanding the privacy concerns and behaviors that first arose out of an emergency, WFH arrangements should now be considered a key part of creating healthy work environments for workers. To do that, we must first understand the privacy concerns and behaviors of those whose “new normal” includes WFH on a regular basis.
Our study aims to provide a detailed account of the privacy concerns workers face and the protective measures they take to prevent them. In particular, this study aims to answer the following main questions: How prevalent is experiencing privacy-invasive scenarios among WFH workers? To what extent does experiencing these scenarios cause discomfort? Does discomfort escalate to cause harm to workers or others? What type of harm do workers experience? What types of scenarios are associated with the highest reported discomfort? What protective measures do some workers take? Why do some workers choose not to take protective measures?
To this end, we surveyed 214 US-based workers who currently work from home regularly about their experiences with privacy-invasive scenarios pertaining to: audio, video, data [^1] For brevity, throughout the paper, we use the word “data” to denote “computer/browsing data.”, and autonomy that can occur in a WFH setting.
@@
    Privacy-invasive scenarios are commonly experienced and caused discomfort to most of our participants. Almost all participants (93.9%) experienced at least one of the privacy-invasive scenarios, and more than half of the participants (65.4%) felt uncomfortable during at least one of the scenarios they experienced.
@@
    Most of the participants who experienced a privacy-invasive scenario that made them feel uncomfortable said that the incident did not escalate to cause harm to them or others (74.3%).
@@
    While scenarios that restrict worker autonomy, such as prohibiting them from turning off the camera or microphone, are not among the most commonly experienced scenarios, they are associated with the highest reported discomfort. Participants often violated or would violate the employer’s autonomy-restricting rules to protect their privacy.
@@
    Participants reported experiencing more discomfort with audio privacy invasions than with video privacy invasions (46.2% versus 28.8% of video privacy-invasive scenarios).
@@
    Smart measures (software/hardware settings or features to prevent privacy-invasive scenarios, e.g. virtual background and noise cancelation features) are not widely adopted compared to manual measures (e.g. manually covering the camera and turning on/off the microphone). When used, more participants reported using settings or features to prevent video scenarios (24.0%) than audio (13.11%), data (8.4%), and autonomy (4.2%) scenarios.
Our findings provide useful insights to organizations’ policymakers and technology designers to identify areas of improvement to provide healthier work environments. In addition, we provide several open questions that can inspire future work.
2. Background and Related Work
2.1. The Shift to WFH
The concept of remote work is not new (McAlpine, 2018; Hayes et al., 2021; Chiru, 2017). It has gained popularity in the past 40 years as computers, smartphones, Internet connectivity, and collaboration platforms were becoming ubiquitous (Chiru, 2017). Before 2020, WFH was often provided voluntarily, as a form of benefit such as reduced commuting (Hayes et al., 2021). However, with the advent of the global COVID-19 pandemic in early 2020, many people who held positions that required them to go to a traditional office space regularly suddenly found themselves required (by their employer or lockdown laws) to stay home, and WFH was no longer a benefit, but a mandate.
WFH mandate created disparate experiences for individuals (Ford et al., 2021). Ford et al. study on software developers’ experiences with WFH found that some developers reported being more productive, while others faced significant challenges (Ford et al., 2021). WFH can impose challenges on some people, such as affording suitable space at home, and the overlap between work-life and home-life, which in turn may affect worker performance and well-being (Fukumura et al., 2021). There are also legal considerations about WFH practices, privacy, and how the two might impact an individual’s livelihood. For example, in the US, the default employment standard is “employment-at-will,” which means that individuals are not guaranteed work and their employers may choose to terminate employment at any time (Varner and Schmidt, 2022). As a result, workers may feel the need to self-censor their home environments to help align with what they believe are the social expectations of others in their organization to preserve their reputation and employment status (Varner and Schmidt, 2022). Thus, privacy-protecting behaviors may also be employment-protecting behaviors.
After COVID-19, many individuals and organizations found that the benefits of WFH outweigh the drawbacks. Many organizations have changed their policies to allow WFH on an ongoing basis (Smite et al., 2023; Christian, 2023; Lufkin, 2022). Thus, understanding the privacy concerns workers face is a key part of creating healthy work environments. Our work is a step in this direction.
2.2. Technologies that Enabled WFH
Technologies that allow workers to work remotely, such as video conferencing, unified communication, and productivity monitoring tools, existed well before the onset of the pandemic and were used regularly by those who had WFH arrangements with their employer (Rysavy and Michalak, 2020). However, the government-issued mandatory COVID-19 lockdown created a need for employers to quickly switch to a new way of working for their entire organization (Hodder, 2020). Video conferencing technologies, such as Microsoft Teams (Microsoft, 2025), Google Meet (Google, 2025), and Zoom (Inc., 2025), became part of workers’ daily lives, and mobile phones (often the worker’s personal property) became an ad hoc office phone. Organizations that had unified communication systems leveraged them in new ways to help workers stay connected (Rysavy and Michalak, 2020).
As COVID-19 lockdown extended from weeks to months of remote work, employers became increasingly interested in ensuring that their workers were productive, which resulted in a sharp increase in the demand for remote monitoring technologies (Mosendz and Melin, 2020). Remote monitoring tools can provide employers with a variety of surveillance functions. These tools may include features ranging from basic time clock functionality (through which the worker can check in and out of their work shift) to remote monitoring tools which are installed on computers to track the time spent on certain applications, documents opened, email read, and keystrokes (Nguyen, 2020; Vatcha, 2020). Some remote monitoring tools can be configured to keep a log of all websites a worker visits and take a screenshot of the worker’s computer screen at regular intervals (Allyn, 2020; Vatcha, 2020). Furthermore, in addition to monitoring worker productivity based on their interaction with their computer, it is not uncommon for employers to opt to use video conferencing tools as a worker surveillance tool, requiring workers to keep their cameras and microphones on throughout their workday (Mosendz and Melin, 2020). Vitak et al. showed that WFH workers have concerns about worker monitoring practices when data collection exceeds the norms and when monitoring limits workers’ autonomy (Vitak and Zimmer, 2023).
Our work complements existing work in this realm and adds new insights to what is known about workers’ autonomy-invading scenarios and workers’ attitudes towards them.
2.3. WFH and Security & Privacy Issues
WFH security and privacy issues are often interrelated (a security attack can cause a privacy breach). Sometimes both types of issues are combined as cybersecurity issues. However, for the sake of clarity, in what follows, we dissect previous work into two categories: security-focused and privacy-focused studies. Our work is considered a privacy-focused study from the human aspect.
2.3.1. Security-Focused Studies
Several studies investigated the security attacks that emerged or increased during the COVID-19 lockdown when WFH was mandated. Pranggono and Arabo provided a summary of notable cybersecurity attacks that increased during the COVID-19 period (Pranggono and Arabo, 2021). Kotak et al. evaluated the various technologies that allow remote work, and introduced a taxonomy of the security threats that these technologies and WFH introduced to organizations (Kotak et al., 2023). Lallie et al. highlighted key cybercrime incidents with respect to the timeline of key global events during the COVID-19 pandemic (Lallie et al., 2021). Other studies provided empirical analysis on particular types of attacks, such as Zoombombing (Ling et al., 2021), phishing (Akdemir and Yenal, 2021; Hoheisel et al., 2023), and ransomware (Lang et al., 2023), to list a few.
Another line of research on security issues and WFH was focused on workers’ and organizations’ perspectives. Georgiadou et al. investigated workers’ security readiness for WFH during the COVID-19 lockdown (Georgiadou et al., 2022). Their findings showed a lack of security readiness, especially in terms of the human aspect. For example, 53% of the workers had not received security guidelines regarding WFH from their organizations, which looks even worse knowing that 44% of them had never experienced WFH (Georgiadou et al., 2022). Alromaih et al. conducted a qualitative study on 20 WFH workers investigating the phenomenon “shadow security behavior,” where workers apply security practices that do not comply with the organization’s security policy (Alromaih et al., 2024). They introduced a model for “personal security” and identified how it interacts with an organization’s security model in the WFH setting (Alromaih et al., 2024). Bispham et al. interviewed six cybersecurity experts on the cybersecurity issues organizations face in the WFH setting during COVID-19 (Bispham et al., 2022). The research found that while WFH raised new cybersecurity challenges, experts indicated that cybersecurity issues did not undermine effective WFH (Bispham et al., 2022). The researchers suggested the need for surveys to gain an empirical understanding of cybersecurity issues during and after COVID-19 WFH setting (Bispham et al., 2022), which is the contribution of our work.
2.3.2. Privacy-Focused Studies
A few recent studies focused on privacy issues in WFH setting. Emami-Naeini et al. conducted a survey on privacy concerns and attitudes toward remote communications in work, socializing, and learning contexts (Emami-Naeini et al., 2021). They found that privacy issues in remote communications are among the issues people care about and affect their comfort in using conferencing tools (Emami-Naeini et al., 2021). They also found that most of the participants felt that they lacked autonomy to choose the conference tool and/or to use the microphone and camera in their remote communications (Emami-Naeini et al., 2021). Prange et al. investigated privacy breaches that occurred as a result of video meetings (Prange et al., 2022). Of the numerous examples of audio and video-based privacy breaches reported by their participants, more than half of them occurred in business-related scenarios, and many related to video conferencing tools broadcasting accidental audio or video sharing of private situations (Prange et al., 2022). Cheetham et al. interviewed 18 participants to identify how the COVID-19 pandemic impacted people’s privacy (Cheetham and Michalec, 2024). Their results provided high-level themes and showed that workers face accidental or inappropriate disclosure, feel they have limited control over their privacy, and are reluctant to discuss privacy with their employers due to concerns about privacy stigma (Cheetham and Michalec, 2024). Whitty et al. conducted a study on the human factor in the WFH environment that includes workers’ cybersecurity training and the incidents they may face (Whitty et al., 2024). While the study is not dedicated to privacy and covered both security and privacy issues, it shed light on a few privacy-invasive incidents such as the unexpected appearance of children at work-related meetings (Whitty et al., 2024). Herder and Gullit examined workers’ privacy-protective behavior with respect to “power distance,” a term used to describe the extent to which less powerful members of an organization accept variations in power distribution (Herder and Gullit, 2022). They found that the presence of a superior did not affect participants’ reported privacy-protective measures in group of peers meetings as much as group size and familiarity with group members did (Herder and Gullit, 2022).
Several privacy-focused studies focused primarily on remote education context. Li et al. and Castelli et al. investigated students’ unwillingness to share their cameras in remote classes (Li et al., 2022; Castelli and Sarvary, 2021). Both studies identified privacy as one of the reasons that led students not to share their cameras in remote classes (Li et al., 2022; Castelli and Sarvary, 2021). Students indicated concerns about their appearance, the appearance of their physical space, or the people around them (Li et al., 2022; Castelli and Sarvary, 2021). Cohney et al. surveyed university instructors and administrators, and analyzed the security of a set of the most popular virtual learning platforms (Cohney et al., 2021). While the study is not solely privacy-focused, the instructor survey shed light on some privacy issues. It showed that nearly half of the instructors have security and privacy concerns, care about recording controls, and some were dissatisfied with the choice of the platform and its configuration by their university (Cohney et al., 2021).
@@ 2.3.2. Privacy-Focused Studies
Technologies designed to protect workers’ privacy can fail under certain assumptions. Hilgefort et al. showed that software features that can be used to enhance privacy while using video conferencing tools, in particular, virtual backgrounds (which can be used to obfuscate surroundings) are vulnerable to attack (Hilgefort et al., 2021). They found that artificial intelligence can be used to effectively remove the blurred virtual backgrounds native to many video conferencing tools, revealing the participants’ true physical setting (Hilgefort et al., 2021).
While themes related to some types of privacy-invasive scenarios were identified in prior work, our scenario-based survey combined a list of common privacy-invasive scenarios from different categories (audio, video, data, and autonomy), providing a more comprehensive approach that allowed us to identify nuanced and novel insights that none of the previous work identified. For example, previous work identified autonomy-invading scenarios; however, our results provide insight on their prevalence and the reported discomfort associated with them compared to other scenario categories. While previous work surfaced workers’ privacy concerns when working from home, our results evaluated to what extent the discomfort resulting from experiencing privacy-invasive scenarios escalates to cause harm to the workers or others, and what types of harm they cause. Such nuanced insights were missing in previous work, although they are essential for informed decisions about WFH policies and arrangements by employers and policymakers to provide healthier work environments. Unlike the samples surveyed in previous work, which included workers and non-workers or focused on the education context (students and/or educators), our sample includes only WFH workers, providing us with deeper insights into the issues workers face while working from home.
@@ 3. Method
We conducted a 214-participant online survey. We obtained ethical approval for the study from the CMU Institutional Review Board (IRB). All participants were presented with an online consent form at the beginning of the survey.
@@ 3.1. Design
Our survey consisted of two parts: a screening survey and a main survey. We used the screening survey to select participants who were currently professional workers (at the time of the study) and working from home regularly (one or more days per week, either in full or hybrid mode), in addition to other screening criteria (e.g. working from home for at least three months), which we detailed in the demographics section of the results.
Only those who met our screening criteria were invited to take the main survey and were immediately redirected to the main survey if they agreed. The screening survey is provided in Appendix A.
To design the main survey questions, three members of the research team developed a set of 14 privacy-invasive scenarios that can occur in a WFH setting, listed in Table 1. Some of these scenarios were inspired by the results of an IRB-approved unpublished 50-participant exploratory survey on WFH privacy behaviors and concerns conducted by some of this paper’s authors and others in 2023. The 14 scenarios were divided into four categories: audio, video, data, and autonomy. The audio and video categories (five scenarios for each category, totaling 10 scenarios) presented leaked audio/video scenarios through the subject’s microphone/camera from five different actors: the subject, another adult, child, pet, and object (e.g. appliances, books, furniture, artwork). The data category (two scenarios) presented leaked data scenarios through the subject’s shared screen. Finally, the autonomy category (two scenarios) presented restricted subject’s autonomy to turn off the microphone/camera. All scenarios were presented in the context of a work-related remote call or meeting from home.
Table 1. The 14 scenarios presented to participants were divided into categories. Scenarios that share a large amount of text are listed once in the shaded rows for each scenario category, starting with three dots (…) that represent the variable part of the scenario that changes according to the individual scenario listed below.
@@
To further characterize WFH scenarios that cause discomfort, after all scenario questions were answered, we presented participants with a list of all the scenarios that they indicated made them feel uncomfortable. Then, we asked participants to select the most memorable scenario and then asked them follow-up questions about the one selected scenario. The follow-up questions included an open-ended question to describe the incident in more detail, and several multiple-choice questions including whether the discomfort escalated to cause harm to themselves or others. If the participant did not report any scenario that made them feel uncomfortable, we asked them to describe a scenario that would make them feel uncomfortable if they were to experience it while working from home and why. The survey concluded with demographic questions.
To avoid security and privacy priming, we did not mention the word privacy or security to participants at any point before or during the survey. The survey was conducted using Qualtrics (Qualtrics, 2024), an online survey platform. The complete screening and main survey questions are included in Appendix A.